. A poster must have 32 sq.in. of printed materials with margins of 4 inches each at the top and 2 inches at each side. Find the dimensions of the whole poster if its area is maximum.


1
Expert's answer
2022-06-13T23:27:15-0400

Let "x=" the length of printed matter, "y="the width of printed matter.

Given "xy=32."

The length of a poster will be "x+2(2)," the width of a poster will be "y+4."

The area of a poster is "A=(x+4)(y+4)."



"xy=32=>y=\dfrac{32}{x}"

Substitute



"A=A(x)=(x+4)(\dfrac{32}{x}+4), x>0""A(x)=32+4x+\dfrac{128}{x}+16"

Find the derivative with respect to "x"



"A'(x)=(48+4x+\dfrac{128}{x})'=4-\dfrac{128}{x^2}"

Find the critical number(s)



"A'(x)=0=>4-\dfrac{128}{x^2}=0""x^2=32""x_1=-4\sqrt{2}, x_2=4\sqrt{2}"

If "0<x<4\sqrt{2}, A'(x)>0, A(x)" increases.


If "x>4\sqrt{2}, A'(x)<0, A(x)" decreases.


The function "A(x)" has a local maximum at "x=4\sqrt{2}."

Since the function "A(x)" has the only extremum for "x>0," then the function "A(x)" has the absolute maximum at "x=4\sqrt{2}" for "x>0."



"y=\dfrac{32}{4\sqrt{2}}=4\sqrt{2}"

The poster is "(4\sqrt{2}+4)\ inches\times (4\sqrt{2}+4)\ inches."